Circular No. 6886

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bn IN NGC 4462 W. D. Li, M. Modjaz, R. R. Treffers, and A. V. Filippenko, University

of California at Berkeley, report the discovery of an apparent supernova during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope. SN 1998bn was found and confirmed on unfiltered CCD images obtained on Apr. 17.3 (mag ∼ 17.4) and 27.3 UT (mag ∼ 13.7), respectively. The object is located at α = 12^{h}29^{m}18^{s}.88, δ = −23^{o}09′49′′.2 (equinox 2000.0), which is ∼ 32′′.3 west and 9′′.7 north of the nucleus of NGC 4462. The new star already appeared in the template image obtained on Apr. 17.3, and brightened ∼ 3.7 mag in 10 days. The Digital Sky Survey does not show any star at the position of SN 1998bn.

SAX J1808.4–3658 = XTE J1808–369 A. B. Giles, K. M. Hill, and J. G. Greenhill, Physics Department, Uni-

versity of Tasmania, report: “We have CCD observations of SAX J1808.4– 3658 that appear to confirm the optical candidate proposed by Roche et al. on IAUC 6885. A set of 23 V -band images covering the ∼ 2-hr x-ray binary period of Chakrabarty and Morgan (IAUC 6877) was obtained with the Mt. Canopus 1-m telescope during Apr. 22.70–23.79 UT. There is a roughly-sinusoidal modulation of 0.12 mag peak-to-peak, with a maximum at ∼ Apr. 22.75. A set of five V images spanning ∼ 0.5 hr obtained on Apr. 18.70 shows the object to be brighter by ∼ 0.1 mag, assuming that the sine modulation is persistent. The intensity decrease between the two data sets provides further evidence for the proposed identification.”

COMET 1998 H1 A. Miyamoto and T. Oribe, Saji Observatory, report (via S. Nakano,

Sumoto) a coma diameter of 12′′ and a tail 3′ long in p.a. 190^{o} on CCD images obtained on Apr. 26.774 UT with a 1.03-m reflector.

Total visual magnitude and coma-diameter estimates: Apr. 26.97 UT, 10.6, 3′ (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 26.99, 10.2, 6′ (M. Meyer, Frauenstein, Germany, 20×100 binoculars).

Corrigenda to IAUC 6883. Line 13, for 23.016 read 23.16. Fourth line from the bottom, for p.a. 155^{o} read p.a. 205^{o}.

1998 April 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green